### v1

Abstract
Do social networks and peer influence shape major life decisions in highly polarized settings? We explore this question by examining how peers influenced the allegiances of West Point cadets during the American Civil War. Leveraging quasi-random variations in the proportion of cadets from Free States, we analyze how these differences affected decisions about which army to join. We find that a higher proportion of classmates from Free States significantly increased the likelihood that cadets from Slave States joined the Union Army, while almost all cadets from Free States joined the Union Army (if they decided to join the war). We further examine how cadets’ decisions affected their military rank and career outcomes. Our findings highlight that peers still influence choices even when they are life-altering and occur during periods of extreme polarization.
JEL Codes: D85, J24, N31, N41
“… for Civil War soldiers, the group cohesion and peer pressure that were powerful factors in combat motivation were not unrelated to the complex mixture of patriotism, ideology, concept of duty, honor and manhood, and community or peer pressure that prompted them to enlist in the first place.”
James M. McPherson, For cause and comrades: Why men fought in the Civil War (1997)
1 Introduction
In a highly polarized world, ideological divides not only reflect personal opinions but also actively shape critical decisions, from health choices like vaccination uptake to alignment in global conflicts. Because of these divides, individuals often grapple with challenging decisions and conflicting incentives. For example, during the COVID-19 pandemic, focus groups and surveys suggest that social expectations—–sometimes at odds with personal preferences—–played a significant role in vaccination decisions (Lin et al., 2022). Similarly, in politically polarized settings, it has been suggested that decisions about voting or expressing political views can be deeply influenced by social networks, as individuals may suppress their private preferences to conform to publicly dominant views (Kuran, 1998). Theory has long suggested that social networks play a critical role in shaping these decisions, not only through peer pressure for conformity or allegiance, but also by providing insight and information and sharing relevant experiences, and providing comradery (e.g., Bernheim, 1994, Jackson, 2008).
Although there is extensive evidence of peer effects in a variety of settings (Benhabib, Bisin and Jackson, 2010; Brechwald and Prinstein, 2011; McGloin and Thomas, 2019; Giletta et al., 2021), it is unclear whether peers still sway decisions when other economic and social forces become enormous and a single decision can be life-altering. Providing causal evidence on how social networks influence behavior in highly polarized and life-changing environments remains challenging because individuals tend to associate with like-minded peers, making it difficult to disentangle the influence of social networks from preexisting shared beliefs and communal forces.
In this paper, we address this challenge by examining the historical context of the American Civil War, one of the most polarized periods in U.S. history. Specifically, we investigate how West Point cadets–—a group that later served as leaders in both the Union and Confederate armies–—were influenced by their peers when deciding their allegiances. Although West Point admissions broadly reflected the geographic distribution of congressional districts over time, annual recruitment patterns varied significantly due to the decentralized nomination process and qualification exams (e.g., Park, 1840). We show that this year-to-year variation in peer composition was quasi-random. Exploiting this feature, we provide evidence of the significant role that peers played in shaping decisions during a moment of extreme polarization, and where other background communal and geographic forces might have seemed otherwise overwhelming.
The American Civil War was a pivotal event in U.S. history, notable not only for its significant political and social impacts but also for the personal conflicts it caused among individuals torn between nationalism and sectionalism. West Point cadets, who later became leaders of both the Union and Confederate armies, were among those who faced split allegiances. Although these cadets played vital roles in the war, there has been limited empirical investigation into the factors influencing their decision on which side to fight for. [^1] There exists a rich historical literature on West Pointers in the Civil War (e.g., Patterson, 2002). A recent political science study, White (2024) documents the importance of economic interest in influencing cadets’ choices. See additional reference below.
West Point, as the premier military training institution of the time, offered a unique environment where cadets from both Northern (Free) and Southern (Slave) States lived, trained, and formed close-knit cohorts prior to the war. This setting makes peer influence a particularly relevant dynamic for study. There is research suggesting that peer influence and “group cohesion” played important roles in motivating people to enlist (and fight), but much of the evidence is anecdotal and drawn from letters and diaries (e.g., see McPherson (1997, 2003); Siebold (2007)). How interactions between cadets from these different regions influenced allegiance during the Civil War has yet to be explored with a careful statistical analysis. Did exposure to northern peers sway southern cadets toward the Union cause, or did the sociopolitical context of their home states dominate and insulate them from peer influence?
We study these questions by digitizing detailed biographical data on over 1,600 West Point cadets from 1820 to 1860 and leveraging quasi-random variations in peer composition. We classify states as slave or free based on whether slavery was legal in 1860 (which corresponds to whether the slave population in the 1860 census exceeded 1%), and we revisit the results with alternative thresholds for robustness. The share of peers from Free States fluctuated annually without clear trends, providing variation in peer exposure (see Section 3.1 for institutional reasons that drive the variation). This variation is not systematically correlated with individual characteristics or home-state differences—whether considering the full sample or restricting to cadets who later participated in the war—consistent with the interpretation of peer composition as quasi-random. This quasi-randomness allows us to overcome common endogeneity concerns in identifying peer effects (e.g., Manski, 1993; Angrist, 2014). In the main text, we focus on the wartime choices of cadets who served in the conflict. The Supplemental Appendix presents additional robustness checks using the full cadet sample.
Our main finding is that a higher proportion of classmates from Free States significantly increased the likelihood that cadets from Slave States joined the Union Army. The effect is sizable: a one-standard deviation increase in the proportion of peers from Free States raised the probability of joining the Union by four percentage points, or 11% of the mean. In contrast, nearly all cadets from Free States joined the Union regardless of cohort composition. The difference suggests a conflict between nationalism and sectionalism faced by cadets from Slave States, who were more susceptible to peer influence. Furthermore, within the Slave States, we find that peer influence mattered little for cadets from states in which the slave population was high (e.g., above 1/3), but mattered strongly for cadets from Slave States with slave populations below 1/3, suggesting an interaction between peer influence and economic considerations rooted in the slave economy. Indeed, the peer influence was large and almost entirely concentrated among cadets from Border and Mid-level Slave States where support for secession was more divided.
We present several additional analyses to examine heterogeneity and test the robustness of our results. Notably, peer effects are stronger when Slave States are defined using thresholds of 5% or 10% slave population, suggesting that states with minimal slave populations resemble Free States in their influence. Moreover, peer influence is more pronounced within the same cohort than across cohorts, indicating the importance of direct, contemporaneous interactions. Similarly, cadets who completed their training exerted greater influence than those who did not.
Because many West Point cadets in our study later emerged as significant military figures, we further examine how their wartime choices influenced their post-war trajectories. Employing both OLS and an instrumental variable strategy (where home state and peer exposure predict allegiance), we find that for those from Slave States, those who sided with the Union experienced lower military ranks in 1865, a diminished chance of becoming generals, and a decreased likelihood of wartime mortality.
The variation in cohort composition used in our main analysis does not pinpoint the mechanism through which peer influence operates. To probe this further, we examine variation in cadets’ likelihood of having co-fought in the Mexican-American War (1846–1848). We find that peer effects were stronger when slave-state cadets had more free-state peers with whom they had shared combat experience, and weaker when their co-combatants were also from Slave States. Due to the limited sample—restricted to slave-state cadets who graduated before 1846—the interaction estimates are sizable but imprecise. While not conclusive, this pattern suggests that shared formative experiences, such as fighting side-by-side, played a role in peer influence.
Our study contributes to the broad social science literature on peer effects (see Benhabib, Bisin and Jackson (2010); Brechwald and Prinstein (2011); McGloin and Thomas (2019); Giletta et al. (2021) for some overviews), by demonstrating how peer influence can shape major life decisions, particularly during moments of extreme polarization. The experiences of West Point cadets during the American Civil War mirror other historical contexts where military academies produced leaders for both sides of major conflicts in divided nations. For instance, the Whampoa Military Academy in the 1920s produced commanders who fought on opposing sides during the Chinese Civil War (e.g., Bai, Jia and Wang, 2023), and the Heroic Military Academy in Mexico trained leaders who both supported and opposed the regime during the Mexican Revolution.
As mentioned above, such decisions are not unique to military cadets; similar choices are faced by ordinary citizens in many divided societies. However, analyzing peer influence in these broader contexts is often more challenging. The distinctive setting of the American Civil War, where allegiances were closely tied to geography, provides a unique opportunity to examine peer influence with a level of identification that is difficult to achieve in other cases.
Our findings also add to the vast body of historical and social sciences literature that examines loyalty decisions during American Civil War. McPherson (1997) draws on soldiers’ letters and diaries to investigate soldiers’ motivations. Analyzing Union Army company data, Costa and Kahn (2003, 2008) emphasize the importance of loyalty to fellow soldiers, which sometimes surpassed commitment to the cause when considering decisions to fight, shirk duty, or desert. Our findings suggest that both cause (the allegiance to one’s home state) and comrades (peer influence) matter and that they also interact with each other. Closely related, Hall, Huff and Kuriwaki (2019) document that slave ownership drove enlistment in the Confederacy, whereas White (2024) explores the allegiance of West Pointers, emphasizing economic interest by showing that one’s past employment related to cash crops (i.e., sugar, indigo, rice, tobacco, and cotton) is linked to the likelihood of West Pointers joining the Confederacy. [^2] Using the same variable on cash-crop employment, we find a similar pattern. But with only a small share of candidates having such a history, we do not find a significant interaction effect between this variable and peer influence. Our observation regarding the influence of the slave economy corroborates these findings, though our primary focus remains on peer influence—a critical factor in loyalty decisions during pivotal historical moments.
Beyond this context, a large literature has examined the determinants of conflict participation, highlighting the importance of interests (as reviewed in Blattman and Miguel (2010)) and social forces (reviewed in Cantoni et al. (2024)). [^3] There is also a long-standing literature in sociology recognizing the importance of personal networks in social and political movements (e.g., McAdam, 1986, Diani and McAdam, 2003). We contribute to the previous literature in terms of the variation in peer composition that identifies peer influence, the differing ways that it works depending on background of a cadet’s state, the decisions made by leaders, and the choice between competing ideologies.
Finally, our study contributes to the expanding literature on leaders and leadership (e.g., Jones and Olken, 2005; Acemoglu and Jackson, 2015; Dippel and Heblich, 2021; Cagé et al., 2023; Bai, Jia and Yang, 2023). Understanding how leaders develop their ideas and loyalties is central to grasping important political and economic outcomes. Our findings highlight the significance of peer influence in shaping future leaders’ decisions during crises, demonstrating how peer influence can drive individual choices that may shape the course of history.
2 The Data
We describe how the data are constructed. Full details appear in Supplemental Appendix A.1.
2.1 The Cadets
We focus on cadets who graduated from West Point Military Academy between 1820 and 1860. We manually collected information from the comprehensive Biographical Register of the Officers and Graduates of the United States Military Academy (Cullum, 1891). The records provide details such as the year of graduation, birth year, home state and appointment state, [^4] Home state and appointment state usually overlap. We use home state in our main analysis and present additional robustness checks using appointment state to define peers in the Supplemental Appendix. and academic rank. To ensure comparability of academic performance, we classify cadets into percentiles based on their graduation ranking. The top 10% were assigned a score of 100, the bottom 10% a score of 10, and so on. Additionally, the records include information on wars and battles one participated in and their military ranks, which we use to determine if one joined the Union or Confederate forces during the Civil War. We present examples of these records and our coding procedure in Supplemental Appendix A.1.
Since the Biographical Register lacks some information about eventual military ranks and death for some cadets who joined the Confederacy, we supplement our data with sources including Rebels from West Point (Patterson, 2002), Southern Historical Society Papers (Maury, 1876), and Confederate Military History (Evans, 1899). We also account for cadets who did not graduate from West Point during 1820-1860, using the Register of Graduates and Former Cadets (United States Military Academy, 1971). Additionally, we cross-reference our data with Wikipedia and epitaph data (Find a Grave, 2024) whenever information about a cadet is accessible.
Data on the economic conditions of cadets’ home states are sourced from the 1860 Census (United States Census Bureau, 1860). Our primary focus is on one key variable that serves as a proxy for economic conditions closely tied to the Civil War: the share of the slave population in the total population. Additionally, we examine other economic proxies, including logged farmland value per capita, logged manufactured product value per capita, and the share of manufacturing employment. These variables are typically highly correlated with each other and are absorbed by state fixed effects in our analysis. In additional analysis, we present the correlations between these variables and war allegiance, confirming that the share of the slave population is the most robust economic predictor of war allegiance.
Following White (2024), we code cadets’ employment histories and use involvement in cash-crop occupations as a proxy for material interests, allowing us to compare this measure with the slave population share. In addition, we leverage data on cadets’ participation in the Mexican-American War (1846–1848) to shed light on potential channels through which peer influence may have operated.
2.2 Geography and the Classification of States by High/Low Slave
We are interested in how cadets’ decisions were shaped by the choices of their peers. To address the challenge of endogeneity in peer choices, we use cadets’ home state as a source of exogenous variation. Specifically, we explore the impact of the proportion of cadets from Free States on the allegiance decisions of cadets from both Free and Slave States. [^5] Our approach is consistent with the recommendation in Angrist (2014), where “research designs that manipulate peer characteristics in a manner unrelated to individual characteristics provide the most compelling evidence on the nature of social spillovers.”
There were 11 states that seceded and formed the Confederate States of America by the end of 1861 (Alabama, Arkansas, Florida, Georgia, Louisiana, Mississippi, North Carolina, South Carolina, Tennessee, Texas, Virginia). There were four states (Delaware, Kentucky, Maryland, and Missouri) that were considered Border States [^6] The District of Columbia is generally also added to the border category, although not a state. and that had significant internal splits in their allegiance, and in some cases even parallel governments.
In the main text we categorize states existing at the beginning of the war [^7] West Virginia was split off from Virginia in 1863, becoming part of the Union, but we treat it as part of Virginia as of 1860. into Slave and Free States based on whether slavery was legal in 1860, which corresponds to using a 1% slave population share in 1860 as a threshold. In the Supplemental Appendix, we show that our findings on peer effects remain (and can become even stronger) when using higher thresholds of 5% or 10% of slave populations. Supplemental Appendix Table S1 lists the classification of states using different thresholds.
The territories that were not yet states were under Union government control and none could secede, and none had significant slave populations. [^8] A few had split allegiances, but overall the territories provided relatively small enlistments, and mostly for the Union. There is only one cadet in our sample from these territories, which we exclude from our analysis. We also exclude the few cadets whose hometowns are outside the U.S.
In our data, 964 and 674 cadets came from Free States and Slave States, respectively. Among those from Free States, 540 joined the armies and 91.9% of them joined the Union. Among those from Slave States, 389 joined the armies and 36.2% of them joined the Union. Supplemental Appendix Table S2 presents the summary statistics of the key variables using the baseline threshold.
3 Results
3.1 Descriptive Evidence
Importance of Home State
First, we establish that a cadet’s home state serves as an effective representation of regional norms and preferences, as evidenced by the considerable impact of the state’s slave population percentage on the cadets’ loyalty. Figure 1 illustrates that the proportion of a state’s slave population is strongly and negatively related to the likelihood of joining the Union, regardless of whether we focus on war participants or all cadets.
A: Probability of joining the Union: war participants
B: Probability of joining the Union: all cadets
Figure 1: Home-state slave share and the probability of joining the Union
Note. This figure plots the relationship between the share of the slave population in a state and the likelihood that cadets from that state would join the Union. Panel A concentrates on the states that participated in the war, while Panel B uses all cadets, including those who did not participate in the war. The circle’s size represents the number of West Point cadets hailing from each state.
In general, the slave population serves as a strong predictor for joining the Union compared to the Confederacy, but it has no (significant) correlation with the decision to participate in the war (Supplemental Appendix Figure S2). Thus, we focus on war participants in our analysis.
Relevance of Peers
In Figure 2(a), we illustrate the correlations between the proportion of free-state peers in each cohort and the likelihood of Union affiliation for cadets from both slave and Free States. The figure reveals a distinctly positive relationship for cadets from Slave States, suggesting peer influence when facing a conflict between nationalism and sectionalism. Conversely, nearly all cadets from Free States joined the Union, which aligns with their lower tendency to experience a conflict. [^9] Of the 44 that joined the confederacy, 11 were from New York and 10 from Pennsylvania. The 44 were from slightly earlier cohorts than average and had slightly higher class rank than average.
Figure 2: Relevance of Peers: Motivating Evidence
(a) Free-State peers and allegiances for two groups of cadets
(b) Peer composition over time
Note. Panel (a) shows the relationship between the share of Free-State peers and the probability of joining the Union, depicted separately for war-participating cadets from Free and Slave States. Peer influence appears stronger among the latter. Panel (b) presents the time-series of peer composition, showing substantial variation over short periods of time.
What Drives Variation in Peer Composition?
As Figure 2(b) shows, there was substantial year-to-year variation in the peer composition of West Point cadets, which enables our main identification of peer effects. This variation stems from two key features of the Academy’s admissions process (e.g., Park, 1840).
First, admissions were designed to achieve geographic proportionality: the number of cadets from each state was intended to reflect its congressional representation, and, roughly, its population. The admission of cadets required a nomination from a member of the House of Representatives (which now has expanded to also include senators). The slot constraints (some of which went unfilled while others were over-subscribed) provides for substantial randomness in home-state composition from cohort to cohort. We find that logged state population in 1820 alone can explain 24% of the variation in the number of admitted cadets across states.
Second, despite this long-run proportionality, annual admissions fluctuated considerably. Several factors contributed to this. Not all congressional districts submitted nominations each year—many representatives failed to exercise their appointment privilege or were unable to identify qualified candidates. Even when nominations were made, candidates had to pass West Point’s entrance exams; and if a nominee failed, the slot could go unfilled or be reassigned through an at-large appointment. Furthermore, prior to 1843, presidential discretion added an additional layer of unpredictability to the process. [^10] In Supplemental Appendix Table S3, we examine the correlations between $Peer_{i,t}$ and $X_{i,s}$. Once the time trend is accounted for, there is no strong relationship between personal characteristics and peer composition. Moreover, the balance tests hold whether we restrict the sample to cadets who joined the war or include all cadets. Thus, we do not see any observable systematic pattern in the peer composition from cohort to cohort, and cannot reject the identification assumption.
These features were interrelated. Because of the underlying commitment to proportionality, a higher number of cadets from a given state in one year often implied fewer admissions from that state in subsequent years.
3.2 Statistical Analysis
Motivated by the descriptive evidence, we estimate the impact of peer influence via the following OLS baseline specification:
$Union_{i,s,t}=\beta_{0}+\beta_{1}Peer_{i,t}+\gamma X_{i,s}+\alpha_{s}+t+\epsilon_{i,s,t},$ (1)
where $Union_{i,s,t}$ denotes whether cadet $i$ from state $s$ and cohort $t$ joined the Union. $Peer_{i,t}$ represents the proportion of peers from Free States. $\alpha_{s}$ represents state fixed effects, whereas $X_{i,s}$ includes a range of personal and state characteristics such as the cadet’s age in 1860, their academic rank at West Point and the proportion of the slave population in their home state. Naturally, state characteristics are absorbed once home state fixed effects are accounted for. We incorporate a linear time trend $t$ to account for potential temporal trends in Union enlistment, as there is a slight upward trend when considering our entire time horizon. To account for potential serial correlations, we employ bootstrapped standard errors with 400 resampling iterations. We also estimate the model using standard errors clustered at the cohort level, and the statistical significances are unchanged.
As mentioned, we focus on those who joined the armies in our baseline analysis, which is the main margin of peer influence. In addition, we also employ a multinomial logit specification where we consider three outcomes: joining the Union, joining the Confederacy, and not joining the war.
3.3 The Impact of Peers on Cadets’ Allegiance
Baseline Estimates
Table 1 displays our main baseline results, indicating that peer influence matters for individuals from Slave States. Column (1) illustrates the raw correlation between the proportion of Free-State peers and the likelihood of joining the Union. Column (2) incorporates personal and state characteristics, while Column (3) further includes state fixed effects. To facilitate the interpretation of coefficients, we present the coefficient for a one standard deviation change in peer exposure. As shown, these estimates, approximately 0.04, are consistent across different models, confirming the quasi-randomness of peer composition. The effect is sizable, representing about 11% of the mean of the dependent variable (0.36).111111 This is a conservative estimate. When using alternative thresholds (5% and 10%), we find that peer effect raises the probability of Slave-State cadets joining the union by 17-19% (see Supplemental Appendix Table S4).
Table 1: Peer composition and allegiance choice
Dependent var Join the Union: war participants
Slave States Free States
(1) (2) (3) (4) (5) (6)
Share Free-State peers (sd) 0.062∗∗ 0.045∗∗ 0.040 0.006 0.004 0.002
(0.025) (0.022) (0.023) (0.013) (0.013) (0.014)
Age in 1860 -0.014 -0.002 0.007 0.005
(0.016) (0.016) (0.007) (0.007)
Class rank 0.001 0.001 -0.001 -0.001
(0.001) (0.001) (0.000) (0.000)
Slave pop. share (sd) -0.214∗∗∗ -0.009
(0.018) (0.015)
Cohort -0.014 -0.003 0.009 0.007
(0.015) (0.015) (0.006) (0.007)
State FEs N N Y N N Y
Dependent var. mean 0.362 0.362 0.362 0.919 0.919 0.919
Observations 389 389 389 540 540 540
R-squared 0.017 0.220 0.268 0.000 0.015 0.028
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Columns (1)-(3) focus on cadets from Slave States (i.e., with a slave population higher than 1%) and Columns (4)-(6) on those from Free States (i.e., with a slave population lower than 1%). The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
Among the personal background variables, the share of the slave population in a cadet’s home state is strongly and negatively associated with joining the Union, underscoring the role of economic factors in allegiance decisions. Supplemental Appendix Table S5 presents correlations between the probability of joining the Union and additional state-level economic proxies, including logged farmland value per capita, logged manufactured product value per capita, and the share of manufacturing employment. Notably, incorporating these proxies does not alter our primary findings on peer effects. Furthermore, the share of the slave population emerges as the most robust economic predictor of war allegiance. All these economic variables are effectively accounted for when controlling for state fixed effects.
Additionally, consistent with Figure 2, we find that peer influence is not significant for cadets from Free States (Columns (4)-(6)). Based on these findings, our subsequent analysis focuses on cadets from Slave States.
Peer Effects by Type of State
As we see in the analysis above, there is no peer influence on cadets in the Free States. We can further divide the Slave States into groups by the relative slave populations. It turns out that the peer influence operates entirely on cadets from the middle range of Slave States that have slave populations below one third of the population.
In Table 2, we break states into three groups: those without slaves, those with slaves up to one third of the population, and those with slave populations above one third. We now see that the peer effect comes entirely from this middle group of states, and in fact the estimated coefficient and significance are larger than that from the previous table.
These findings suggest that in High-Slavery States, economic imperatives and entrenched social norms overrode peer influence, muting its effect.
Table 2: Peer composition and allegiance choice: different groups of states
Dependent var Join the Union: war participants
Heavy-Slave States Border and Mid-Slave States    Free States
Slave share >33% 1%-33% $\leq$ 1%
(1) (2) (3)
Share Free-State peers (sd) 0.009 0.062∗∗ -0.005
(0.033) (0.030) (0.014)
Share Border-State peers (sd) -0.027 0.022 -0.021
(0.030) (0.032) (0.017)
Age in 1860 -0.056∗∗ 0.025 0.007
(0.023) (0.024) (0.007)
Class rank 0.002 $0$0.001
(0.001) (0.001) (0.000)
Cohort -0.055∗∗ 0.024 0.008
(0.022) (0.023) (0.007)
State FEs Y Y Y
Dependent var.mean 0.115 0.476 0.919
Observations 122 267 540
R-squared 0.168 0.190 0.032
Note. The sample for this analysis consists of all war-participating cadets. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
Relevance of Co-fighting in the Mexican-American War
A significant share—34.6%—of cadets in our dataset participated in the Mexican-American War (1846–1848), offering a valuable context to examine how shared experiences shape peer influence. For each cadet, we compute the proportion of peers who also fought in the war, distinguishing between those from slave and Free States.
If shared wartime experience intensifies peer effects, we would expect our core finding on peer influence to be stronger when a cadet had more co-fighting peers from Free States, and weaker when more such peers were from Slave States. This is precisely what we observe (see Supplemental Appendix Table S6), focusing cohorts between 1820 and 1845. Although the limited sample size prevents precise estimation of this interaction effect, the magnitude is economically meaningful, suggesting that shared combat experience enhances the intensity of peer influence.
3.4 Additional Analyses
We conduct six supplementary analyses to explore (1) whether peer influence within a shared cohort is of greatest significance (compared to adjacent cohorts), (2) how the effect of peers on cadets varied across different various time-frames, (3) potential differences between peers who graduated and those who dropped out, (4) how incorporating non-participants of the war affects our conclusions, (5) how accounting for cash-crop employment shapes peer influence, and (6) estimates based on defining peers by appointment state rather than home state. We describe the results from these analyses below.
Influence by Cohort
We analyze peers by grouping them into cohorts and display the coefficients for peers within the same cohort, as well as for cohorts from t-1 to t-2 and t+1 to t+2, in Supplemental Appendix Figure S3. The coefficients in the figure are obtained from a single regression that includes all controls from our baseline analysis. The result reveals that peer influence is most important within the same cohort. In contrast, although there is some degree of influence from earlier cohorts, future cohorts do not show comparable levels of influence.
Peer Influence pre-post 1850
Although the issue of slavery was highly contentious throught the period we study, it intensified in the 1850sand nearly led to the collapse of the United States. In an effort to avert civil war, Congress passed the Compromise of 1850, which included the Fugitive Slave Law that obligated law enforcement nationwide to assist in capturing alleged runaway slaves, and many in the north resisted its enforcement (McPherson, 2003). Against this backdrop, we use 1850 as a dividing point to assess whether peer influence experienced notable changes before and after that year. As illustrated in Supplemental Appendix Table S7, peer influence (though estimated imprecisely due to limited sample sizes) persisted over time, with a similar effects on either side of this divide.
Graduate vs. Dropout Peer Influence
We focus on graduate peer influence in our main analysis as they spent four years together at West Point. In Supplemental Appendix Table S8, we consider dropouts and calculate peer influence based on them as a comparison. Our estimates are robust to including dropouts in our calculation of peers. However, when separating graduating peers from dropout peers, we find that it is the graduating peers rather than the dropout peers that drives our the peer influence.
Including Those Who Were Not in the Armies
As discussed above, peer composition is not systematically correlated with one’s choice of participating in the war or not. Thus, we have focused on joining the Union versus the Confederacy for simplicity. If we further employ a multinomial logit model to consider whether a cadet joined the Union, joined the Confederacy or did not participate in the war, we again find that peer composition is not predictive for whether a cadet participated in the war, while our main finding of peer influence on joining the Union vs Confederacy holds. These results are presented in Supplemental Appendix Table S9. According to these estimates on the log of the odds ratios, a one-standard deviation increase in Free-State peers increased probability of joining the Union by approximately 4.4 percentage points, higher than our baseline estimates.
Considering Cash-Crop Employment
White (2024) finds that a cadet’s history of employment in cash-crop agriculture (“the graduate was recorded as having spent time as a “planter” (a plantation owner) or held any job related to the production, processing, sale, or export of cotton, indigo, rice, sugar, or tobacco”) is negatively associated with the likelihood of joining the Union. That relationship that holds in our data as well (Supplemental Appendix Table S10). Our primary interest, however, is in whether peer influence is moderated by this employment background.
Unlike the results based on home-state slave share, we find no clear interaction between peer influence and cash-crop employment history (Supplemental Appendix Table S11). This may be due in part to the relatively small proportion—only 7%—of cadets from Slave States having such employment experience.
Considering Appointment State
We have seen that a cadet’s home state serves as an important reference point for allegiance norms. We also examine the appointment state (which heavily overlaps with the home state) and calculate peer composition using the same method as in our main analysis. The analysis is robust to this alternative, as seen in Supplemental Appendix Table S12, we obtain a similar comparable peer effect, as appointment state and home state were often the same.
3.5 Career Outcomes
We know the post-war outcomes of many who joined an army, allowing us to investigate how their decisions during this critical historical juncture influenced their life results, such as military rank and survival probability.
As indicated in Columns (1)-(2) of Table 3, joining the Union correlates with a lower military rank in 1865 and a reduced probability of attaining a rank of General-level positions. This aligns with the fact that the Confederate army, starting from scratch, enabled quicker promotions and a higher likelihood of becoming a general for people with officer training. It also had a higher ratio of generals to enlisted mean. On the other hand, joining the Union is associated with a lower risk of dying in the Civil War, reflecting the higher fatality rate on the Confederacy side, as reported in Column (3).
Motivated our earlier analysis, we employ peer exposure and home state as instrumental variables for joining the Union. The magnitudes of the IV estimate resemble the OLS estimates, as shown in Columns (4)-(6). As previously mentioned, cadets from states with a large slave population (more than one third) did not respond significantly to peer influence. Omitting those cadets deliver similar estimates, despite the reduction in sample size, as seen in Panel B.
Moreover, in Supplemental Appendix Table S13, we investigate the correlation between academic rank and military rank in both the Union and Confederate armies. We find that West Point performance was more predictive of outcomes for the Union Army and less so for the Confederacy.
Table 3: Subsequent outcomes in 1865
OLS IV
Rank General Died Rank General Died
(1) (2) (3) (4) (5) (6)
A. Slave-State cadets (>1%)
Joining the Union -1.556∗∗∗ -0.346∗∗∗ -0.168∗∗∗ -1.952∗∗∗ -0.369∗∗∗ -0.137
(0.201) (0.058) (0.048) (0.351) (0.099) (0.071)
Controls Y Y Y Y Y Y
State FEs Y Y Y Y Y Y
Dependent var. mean 6.974 0.435 0.207 6.974 0.435 0.207
Observations 382 382 389 382 382 389
R-squared 0.314 0.187 0.095
B. Slave-State cadets (1%-33%)
Joining the Union -1.510∗∗∗ -0.318∗∗∗ -0.171∗∗∗ -1.870∗∗∗ -0.373∗∗ -0.161
(0.220) (0.062) (0.051) (0.509) (0.148) (0.093)
Controls Y Y Y Y Y Y
State FEs Y Y Y Y Y Y
Dependent var. mean 6.784 0.402 0.184 6.784 0.402 0.184
Observations 264 264 267 264 264 267
R-squared 0.310 0.187 0.081
Note. The sample for Panel A consists of the war-participating cadets from Slave States (i.e., with a slave population higher than 1%), and the sample for Panel B excludes those from the High-Slave States who were not significantly influenced by peers. Columns (1)-(3) use OLS estimation, while Columns (4)–(6) present IV estimates using birthplace indicators and the share of peers from Free States as instruments for Union affiliation. Columns (1) and (4) feature the military rank as the dependent variable, classified into 11 levels, with General assigned 11 and Third Lieutenant assigned 1. General and Died are binary variables. The rank information on seven cadets is unknown. The standard errors reported in the tables are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
4 Discussion
Our study highlights the significant role of peer exposure in shaping West Point cadets’ allegiance choices during the American Civil War, especially for those from Slave States who faced a tension between nationalism and sectionalism among other tensions. Importantly, we find that the strength of peer influence was modulated by the economic interests of cadets’ home states: in regions where the slave economy was overwhelmingly dominant, the influence of peers was less pronounced. This suggests that when economic stakes were particularly high and slavery more prevalent, cadets were less susceptible to peer-driven shifts in loyalty. This finding is also consistent with the fact that states with heavier slave shares were the first to secede and the level of debate in states was at least partly reflective of slave shares (Crofts, 2014). Thus, the places in which cadets could make choices without seeming disloyal were states with relatively lower slave population shares.
Although our analysis is a case study, it is a useful proof of concept. The decisions that West Point cadets made during the U.S. Civil War parallel those that arise in many divided societies, where individuals must choose between between highly-polarized opposing sides during critical moments. We have shown that in such a context, peer interactions can still play a decisive and significant role.
Our study also provides a road map for identifying such peer effects in other high-stakes and high-polarization settings. This sort of peer influence is challenging to quantify and identify causally through anecdotal evidence alone. It is also not enough to identify correlation in people’s behaviors, given that friendships are endogenous and people are subject to common (potentially unobserved) factors (Angrist, 2014). West Point’s environment—characterized by structured, close-knit interactions among cadets from diverse geographic backgrounds—provided us an opportunity to analyze the influence of quasi-random peer exposure, and its interplay with background allegiances. By using variation in a cadet’s peers’ state-of-origin for identification, we have something that is exogenous to other key things influencing a given cadet’s decision. Similar techniques can be used in other high-stakes contexts.
That said, our ability to definitively trace the mechanisms behind peer influence is limited by the nature of our historical data. Our specification that examines cadets’ choices based on the composition of their cohort by the states of origin of their peers enables us to causally identify peer influence. Nonetheless, it does not provide insight into why peers influence cadets’ decisions. One curious aspect is that the peer influence only affected cadets from states with slaves. This could be due to the fact that West Point was an institution in which cadets were being trained for the U.S. Army, and so only cadets from Slave States faced a conflicting allegiance. Our evidence on how joint participation in the Mexican-American War amplifies peer influence suggests that sharing experiences matters. Theoretically, however, we are still left with the fact that peer effects could have operated in at least three ways: (1) by fostering communication and persuasion, whereby cadets from Slave States may have been convinced to support the Union cause through exposure to pro-Union peers, (2) by building friendship bonds that influenced cadets’ decisions to fight alongside their peers, or (3) by shifting perceptions of the likely outcome of the war, where cadets from Free States could have led their peers to believe that the Union was more likely to prevail. Our data, however, do not allow us to disentangle these, or other mechanisms, and all could have operated together (which would be consistent with some evidence from personal letters and diaries (McPherson, 1997)). Further investigation of peer influence on allegiances in modern contexts, where additional data are more readily available, could offer additional insights into how such mechanisms operate.
Data Availability: all data will be made available upon publication.
Code Availability: all code will be made available upon publication.
Supplemental Appendix for “Comrades and Cause: Peer Influence on West Point Cadets’ Civil War Allegiances,” by Guo, Jackson, Jia
[…]
\@afterheading
Appendix A Data Construction and Description
A.1 Data Construction
We construct a dataset of West Point cadets from 1820 to 1860. We manually collect and organize information on graduates, including their Union or Confederate affiliation during the Civil War, class ranking at graduation, place of birth, year of birth, and military rank in 1865, from the Biographical Register of the Officers and Graduates of the United States Military Academy and other sources. We first present two examples of coding from main source in Figure S1.
A: Example of coding the the Union
B: Example of coding the Confederacy
Figure S1: Text examples
Note. This figure presents text examples from Cullum (15). We use these texts as our main data source and complement the data with information from additional sources
Civil War Affiliation
We determine military allegiance based on whether the individual served with the Union or the Confederacy during the Civil War. This is exemplified by statements such as “He took part in the Rebellion of 1861-66 against the United States” or “He served in the Rebellion of the Seceding States.”
For those whose biographies do not clearly specify a side, we deduce their affiliation from the forces they served with during the war, including volunteer armies and local militias. For instance, service in the New York State Military during the Civil War categorizes them as Union members.
We cross reference our data with additional sources including Find a Grave, Wikipedia, Rebels from West Point, and Southern Historical Society Papers.
Class Rank at Graduation
Since the number of cadets graduating in each West Point cohort varied, direct comparisons of class ranks across different cohorts are not meaningful. To ensure comparability, we classify cadets into percentiles based on their graduation ranking. The top 10% were assigned a score of 100, the bottom 10% a score of 10, and so on.
Year of Birth
In a few cases where an individual’s year of birth was unavailable in the Biographical Register , we supplement this information by using details from Find a Grave.
Military Rank in 1865
For cadets who joined the Union, we determine their official military rank in 1865 from personal biographies. For cadets who joined the Confederacy, we refer to sources including Rebels from West Point, Southern Historical Society Papers, Confederate Military History, and Wikipedia, for military rank descriptions.
If the personal information did not include a specific military rank in 1865, we approximated it based on the individual’s military career trajectory. For example, if an individual’s career record states: {’1846: Second Lieutenant’, ’1864: Major’, ’1870: Colonel’}, we assign the rank held closest to 1865, but no later than 1865. In this case, the individual’s rank in 1865 would be Major. Since every change in a person’s military rank is recorded, if there is no mention of a change, we assume the rank remained the same. We focus on formal military ranks rather than honorary ranks and volunteer military ranks. Military ranks are grouped into 11 levels, with General ranked as 11 and Third Lieutenant ranked as 1.
A.2 Classification of States
Table S1: Classification of States
Slave share (%) 1% threshold 5% threshold 10% threshold 4 groups
South Carolina 57.2 Slave-State $>$5% $>$10% Heavy-Slave State
Mississippi 55.2 Slave-State $>$5% $>$10% Heavy-Slave State
Louisiana 46.9 Slave-State $>$5% $>$10% Heavy-Slave State
Alabama 45.1 Slave-State $>$5% $>$10% Heavy-Slave State
Florida 44.0 Slave-State $>$5% $>$10% Heavy-Slave State
Georgia 43.7 Slave-State $>$5% $>$10% Heavy-Slave State
North Carolina 33.7 Slave-State $>$5% $>$10% Heavy-Slave State
Virginia 30.7 Slave-State $>$5% $>$10% Mid-Slave State
Texas 30.2 Slave-State $>$5% $>$10% Mid-Slave State
Arkansas 25.5 Slave-State $>$5% $>$10% Mid-Slave State
Tennessee 24.8 Slave-State $>$5% $>$10% Mid-Slave State
Kentucky 19.5 Slave-State $>$5% $>$10% Border-State
Maryland 12.7 Slave-State $>$5% $>$10% Border-State
Missouri 9.7 Slave-State $>$5% $<$10% Border-State
District of Columbia 4.4 Slave-State $<$5% $<$10% Border-State
Delaware 1.6 Slave-State $<$5% $<$10% Border-State
New Jersey 0.01 Free-State $<$5% $<$10% Free-State
New York 0.0 Free-State $<$5% $<$10% Free-State
Pennsylvania 0.0 Free-State $<$5% $<$10% Free-State
Ohio 0.0 Free-State $<$5% $<$10% Free-State
Illinois 0.0 Free-State $<$5% $<$10% Free-State
Indiana 0.0 Free-State $<$5% $<$10% Free-State
Massachusetts 0.0 Free-State $<$5% $<$10% Free-State
Wisconsin 0.0 Free-State $<$5% $<$10% Free-State
Michigan 0.0 Free-State $<$5% $<$10% Free-State
Iowa 0.0 Free-State $<$5% $<$10% Free-State
Maine 0.0 Free-State $<$5% $<$10% Free-State
Connecticut 0.0 Free-State $<$5% $<$10% Free-State
California 0.0 Free-State $<$5% $<$10% Free-State
New Hampshire 0.0 Free-State $<$5% $<$10% Free-State
Vermont 0.0 Free-State $<$5% $<$10% Free-State
Rhode Island 0.0 Free-State $<$5% $<$10% Free-State
Minnesota 0.0 Free-State $<$5% $<$10% Free-State
Oregon 0.0 Free-State $<$5% $<$10% Free-State
Note. This table presents different ways we employ to classify home states of West Point cadets.
A.3 Summary Statistics
Table S2: Summary statistics
Mean Std. Dev. Min. Max. Obs.
Panel A: Heavy-Slave States (slave share $>$ 33%)
Joining the Union (war participants) 0.115 0.320 0 1 122
Joining the Union (all) 0.066 0.249 0 1 211
Joining the war 0.578 0.495 0 1 211
Class rank 53.791 30.156 10 100 211
Age in 1860 40.652 11.273 23 65 211
Slave pop. share 45.313 9.410 33.4 57.2 211
Cohort 1841.104 11.636 1820 1860 211
1865 Career rank 5.609 2.601 2 11 207
1865 General 0.295 0.457 0 1 207
Died in civil war 0.180 0.385 0 1 211
Panel B: Border and Mid-Slave States ( 1%$<$slave share $\leq$ 33%)
Joining the Union (war participants) 0.476 0.500 0 1 267
Joining the Union (all) 0.274 0.447 0 1 463
Joining the war 0.577 0.495 0 1 463
Class rank 54.795 28.121 10 100 463
Age in 1860 41.830 10.448 22 62 463
Slave pop. share 20.748 9.427 1.6 30.7 463
Cohort 1839.780 10.870 1820 1860 463
1865 Career rank 5.317 2.399 1 11 460
1865 General 0.241 0.428 0 1 460
Died in civil war 0.117 0.322 0 1 460
Panel C: Free States (slave share $<$ 1%)
Joining the Union (war participants) 0.919 0.274 0 1 540
Joining the Union (all) 0.515 0.500 0 1 964
Joining the war 0.560 0.497 0 1 964
Class rank 56.857 28.725 10 100 964
Age in 1860 41.527 10.931 21 63 964
Slave pop. share 0.000 0.002 0 0.01 964
Cohort 1840.354 11.364 1820 1860 964
1865 Career rank 4.711 2.135 2 10 959
1865 General 0.137 0.344 0 1 959
Died in civil war 0.091 0.288 0 1 956
Note. This table presents summary statistics for the key variables used in our analysis.
A.4 Correlations between Slave Share and Allegiances
A: Probability of joining the Union
B: Probability of joining the Confederacy
C: Probability of being neutral
Figure S2: Slave share and cadet’s choices
Note. The sample for this analysis consists of all cadets, including those who did not participate in the war. This figure illustrates the correlations between the proportion of slaves in a state and the likelihood of a cadet from that state opting to join the Union (A), the Confederacy (B), or abstain from participating (C).
A.5 Peers and Personal Characteristics
Table S3: Correlations between peers and personal characteristics
Dependent var (sd) Share Free-State peers:1820-1860
War participants All
Slave States Free States Slave States Free States
(1) (2) (3) (4) (5) (6) (7) (8)
Age in 1860 -0.007 0.011 -0.011∗∗∗ -0.009 -0.001 -0.003 -0.005 0.008
(0.005) (0.033) (0.004) (0.026) (0.004) (0.029) (0.003) (0.020)
Class rank -0.002 -0.002 0.001 0.001 -0.001 -0.001 0.000 0.000
(0.002) (0.002) (0.001) (0.001) (0.001) (0.001) (0.001) (0.001)
Slave pop. share (sd) -0.097 -0.097 -0.016 -0.015 -0.022 -0.022 -0.048 -0.047
(0.050) (0.051) (0.040) (0.041) (0.040) (0.040) (0.030) (0.030)
Cohort (t) 0.017 0.002 -0.002 0.012
(0.032) (0.026) (0.028) (0.019)
joint test $p$-value 0.124 0.217 0.046 0.093 0.817 0.918 0.101 0.114
Observations 389 389 540 540 674 674 964 964
R-squared 0.016 0.016 0.013 0.013 0.002 0.002 0.006 0.006
Note. This table presents the correlations between peer composition and individual traits. After adjusting for a linear time trend (Cohort (t)), the joint test on overall correlations are statistically undistinguished from zero. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
Appendix B Additional Results
B.1 Alternative Thresholds
Table S4: Peer composition and allegiance choice: different thresholds
Dependent var Join the Union: war participants
Thresholds: 5% Thresholds: 10%
Slave States Free States Slave States Free States
(1) (2) (3) (4) (5) (6) (7) (8)
Share Free-State peers (sd) 0.052∗∗ 0.053∗∗ -$0$0.004 0.059∗∗ 0.058∗∗ -0.002 -0.005
(0.024) (0.024) (0.013) (0.013) (0.023) (0.023) (0.012) (0.013)
Age in 1860 -0.010 -0.001 0.006 0.005 -0.016 -0.006 0.007 0.007
(0.016) (0.016) (0.007) (0.007) (0.018) (0.018) (0.007) (0.007)
Class rank 0.001 0.001 -0.001 -0.001 0.001 0.001 -0.001 -0.001
(0.001) (0.001) (0.000) (0.000) (0.001) (0.001) (0.000) (0.000)
Slave pop. share (sd) -0.179∗∗∗ -0.039∗∗ -0.177∗∗∗ -0.070∗∗∗
(0.019) (0.019) (0.019) (0.020)
Cohort -0.012 -0.002 0.008 0.007 -0.017 -0.007 0.009 0.008
(0.015) (0.016) (0.007) (0.007) (0.017) (0.017) (0.007) (0.007)
State FEs N Y N Y N Y N Y
Dependent var. mean 0.318 0.318 0.910 0.910 0.310 0.310 0.902 0.902
Observations 352 352 577 577 339 339 590 590
R-squared 0.168 0.219 0.035 0.053 0.172 0.224 0.065 0.081
Note. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
B.2 Home State Economic Proxies and War Allegiances
Table S5: Home state economic proxies and war allegiance
Dependent var Join the Union: war participants
Slave States Free States
(1) (2) (3) (4) (5) (6) (7) (8)
Share Free-State peers (sd) 0.045∗∗ 0.045∗∗ 0.043∗∗ 0.042 0.004 0.003 0.003 0.003
(0.021) (0.021) (0.021) (0.022) (0.012) (0.012) (0.011) (0.012)
Age in 1860 -0.014 -0.014 -0.005 -0.004 0.007 0.007 0.007 0.007
(0.016) (0.016) (0.016) (0.016) (0.007) (0.007) (0.007) (0.007)
Class rank 0.001 0.001 0.001 0.001 -0.001 -0.001 -0.001 -0.001
(0.001) (0.001) (0.001) (0.001) (0.000) (0.000) (0.000) (0.000)
Cohort -0.014 -0.015 -0.006 -0.005 0.009 0.009 0.009 0.009
(0.015) (0.015) (0.015) (0.015) (0.007) (0.007) (0.007) (0.007)
Slave pop. share (sd) -0.214∗∗∗ -0.215∗∗∗ -$0∗$114∗∗∗ -0.009 -0.012 -0.013 -0.013
(0.018) (0.019) (0.040) (0.040) (0.014) (0.015) (0.015) (0.015)
ln Farmland value per capita 0.011 0.019 0.034 0.055 0.065 0.056
(0.049) (0.049) (0.049) (0.044) (0.056) (0.065)
ln Manufactured product per capita 0.234∗∗∗ 0.097 0.009 0.021
(0.075) (0.128) (0.032) (0.052)
Manufacturing employment share (sd) 0.074 -0.009
(0.057) (0.033)
Dependent var. mean 0.362 0.362 0.362 0.362 0.919 0.919 0.919 0.919
Observations 389 389 389 389 540 540 540 540
R-squared 0.220 0.220 0.238 0.242 0.015 0.019 0.019 0.020
Note. This table shows the association between war allegiances and several home state economic proxies. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
B.3 Peer influence and shared Mexican-American War experience
Table S6: Peer composition and allegiance choice: Mexican-American War
Dependent var Join the Union: war participants
Slave States
Slave state cadets (>1%) Slave state cadets (1%-33%)
(1) (2) (3) (4) (5) (6)
Share Free-State peers join Mex.-Am. War (sd) $\times$ Share Free-State peers (sd) 0.246∗∗ 0.204 0.208 0.297 0.316 0.328∗∗
(0.121) (0.128) (0.149) (0.164) (0.171) (0.166)
Share Slave-State peers join Mex.-Am. War (sd) $\times$ Share Free-State peers (sd) -0.181 -0.155 -0.155 -0.227 -0.242 -0.254
(0.117) (0.124) (0.141) (0.158) (0.162) (0.156)
Share Free-State peers (sd) 0.029 0.020 0.021 0.028 0.010 0.008
(0.036) (0.033) (0.035) (0.042) (0.045) (0.045)
Share Free-State peers join Mex.-Am. War (sd) -0.152 -0.148 -0.140 -0.268 -0.247 -0.267
(0.125) (0.122) (0.136) (0.147) (0.143) (0.144)
Share Slave-State peers join Mex.-Am. War (sd) 0.129 0.120 0.120 0.237 0.221 0.244
(0.126) (0.121) (0.134) (0.145) (0.143) (0.143)
Controls N Y Y N Y Y
State FEs N N Y N N Y
Dependent var. mean 0.405 0.405 0.405 0.515 0.515 0.515
Observations 185 185 185 136 136 136
R-squared 0.036 0.253 0.294 0.056 0.171 0.224
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Share Free-State peers join Mexican-American War: The proportion of Free-State peers participating in Mexican-American war with themselves. If the cadet did not participate in the Mexican-American War, the proportion is 0. Share Slave-State peers joining the Mexican-American war: The proportion of Slave-State peers participating in Mexican-American War with themselves. If the cadet did not participate in the Mexican-American War, the proportion is 0. Columns (1)-(3) focus on cadets from Slave States. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
B.4 Peer Influence by Cohort
Figure S3: Peer influence by cohort
Note. The sample for this analysis consists of the war-participating cadets from Slave States (i.e., with a slave population higher than 1%). This figure reports the coefficients of peers from the same cohort $t$ and those from three cohorts before and after $t$. The coefficients are derived from a multiple regression that includes all controls from our baseline analysis. The bars represent 90% confidence intervals.
B.5 Peer Influence pre and post 1850
Table S7: Peer composition and allegiance choice before and after 1850
Dependent var Join the Union: war participants
1820-1849 1850-1860
(1) (2) (3) (4)
Share Free-State peers (sd) 0.041 0.034 0.044 0.048
(0.027) (0.029) (0.035) (0.036)
Age in 1860 0.006 0.026 -0.043 -0.035
(0.022) (0.022) (0.026) (0.029)
Class rank 0.001 0.002 $0$0.000
(0.001) (0.001) (0.001) (0.001)
Slave pop. share (sd) -0.224∗∗∗ -0.207∗∗∗
(0.023) (0.031)
Cohort 0.005 0.024 -0.028 -0.020
(0.022) (0.022) (0.026) (0.029)
State FEs N Y N Y
Dependent var.mean 0.391 0.391 0.318 0.318
Observations 235 235 154 154
R-squared 0.227 0.281 0.220 0.333
Note. The sample for this analysis consists of the war-participating cadets from Slave States (i.e., with a slave population higher than 1%). The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
B.6 Graduate Peers vs. Dropout Peers
Table S8: Considering dropout peers
Dependent var Join the Union:war participants
Slave States
(1) (2) (3) (4)
Share Free-State total peers (sd) 0.043 0.039
(0.023) (0.024)
Share Free-State graduates (sd) 0.044 0.039
(0.024) (0.024)
Share Free-State dropouts (sd) 0.001 0.001
(0.021) (0.021)
Age in 1860 -0.013 -$0$0.014 -0.001
(0.016) (0.016) (0.015) (0.016)
Class rank 0.001 0.001 0.001 0.001
(0.001) (0.001) (0.001) (0.001)
Slave pop. share (sd) -0.216∗∗∗ -0.216∗∗∗
(0.019) (0.019)
Cohort -0.015 -0.003 -0.016 -0.003
(0.015) (0.015) (0.015) (0.015)
State FEs N Y N Y
Dependent var.mean 0.362 0.362 0.362 0.362
Observations 389 389 389 389
R-squared 0.219 0.266 0.220 0.267
Note. The sample for this analysis consists of the war-participating cadets from Slave States (i.e., with a slave population higher than 1%). Columns (1)-(2) consider all peers, including those who dropped out. Columns (3)-(4) compare the graduate peers with dropout peers. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
B.7 Multinomial Logit Results
Table S9: Peer influence and allegiance choice: multinomial logit results
Reference group Join the Confederacy
(1) (2) (3)
Share Free-State peers (sd): Join the Union 0.267∗∗ 0.261∗∗ 0.269∗∗
(0.104) (0.115) (0.119)
Share Free-State peers (sd): Not in war 0.044 0.040 0.055
(0.092) (0.111) (0.114)
Controls N Y Y
State FEs N N Y
Dependent var. mean 1.214 1.214 1.214
Observations 674 674 674
Pseudo R-squared 0.005 0.179 0.203
Note. The sample for this analysis consists of all cadets from Slave States, including those who did not join either army. The results show that peer influence mainly affected whether to join the Union or Confederacy rather than whether to engage in the war. The controls are the same as in our previous analyses. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
B.8 Considering Cash-Crop Employment
Table S10: Peer composition and allegiance choice: Cash crops
Dependent var Join the Union: war participants
Slave States Free States
(1) (2) (3) (4) (5) (6)
Share Free-State peers (sd) 0.062∗∗ 0.044∗∗ 0.039 0.006 -$0$0.003
(0.025) (0.022) (0.022) (0.013) (0.012) (0.013)
Age in 1860 -0.014 -0.003 0.007 0.005
(0.015) (0.016) (0.007) (0.007)
Class rank 0.001 0.001 -0.001 -0.001
(0.001) (0.001) (0.000) (0.000)
Slave pop. share (sd) -0.201∗∗∗ -0.005
(0.019) (0.013)
Cash crop employment -0.264∗∗∗ -0.247∗∗∗ -0.624∗∗∗ -0.635∗∗∗
(0.054) (0.052) (0.187) (0.184)
Cohort -0.016 -0.005 0.008 0.007
(0.015) (0.015) (0.006) (0.006)
State FEs N N Y N N Y
Dependent var. mean 0.362 0.362 0.362 0.919 0.919 0.919
Observations 389 389 389 540 540 540
R-squared 0.017 0.236 0.282 0.000 0.081 0.096
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Cash crops is a dummy variable that equals 1 when Graduates with connections to Southern cash crops. Columns (1)-(3) focus on cadets from Slave States and Columns (4)-(6) on those from Free States. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
Table S11: Peer composition and allegiance choice: interaction term of cash crops
Dependent var Join the Union: war participants
Slave States Free States
(1) (2) (3) (4)
Cash crop $\times$ Share Free-State peers (sd) 0.055 0.036 0.022 0.032
(0.057) (0.062) (0.709) (0.683)
Share Free-State peers (sd) 0.042 0.037 -$0$0.003
(0.023) (0.024) (0.012) (0.013)
Age in 1860 -0.015 -0.003 0.007 0.005
(0.016) (0.016) (0.007) (0.007)
Class rank 0.001 0.001 -0.001 -0.001
(0.001) (0.001) (0.000) (0.000)
Slave pop. share (sd) -0.203∗∗∗ -0.005
(0.019) (0.013)
Cash crop employment -0.253∗∗∗ -0.240∗∗∗ -0.613 -0.618
(0.056) (0.053) (0.670) (0.649)
Cohort -0.017 -0.005 0.008 0.006
(0.015) (0.015) (0.006) (0.006)
State FEs N Y N Y
Dependent var. mean 0.362 0.362 0.919 0.919
Observations 389 389 540 540
R-squared 0.237 0.282 0.081 0.096
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Cash crops is a dummy variable that equals 1 when Graduates with connections to Southern cash crops. Columns (1)-(2) focus on cadets from Slave States and Columns (3)-(4) on those from Free States. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
B.9 Considering Appointment State Peers
Table S12: Peer composition and allegiance choice: home state and appointed state
Dependent var Join the Union: war participants
Slave States Free States
(1) (2) (3) (4) (5) (6)
Share Free-State peers (sd) 0.056∗∗ 0.046∗∗ 0.039 -0.003 -0.006 -0.008
Appointed state (0.024) (0.021) (0.022) (0.012) (0.013) (0.013)
Contronls N Y Y N Y Y
State FEs N N Y N N Y
Dependent var. mean 0.362 0.362 0.362 0.919 0.919 0.919
Observations 389 389 389 540 540 540
R-squared 0.014 0.220 0.268 0.000 0.016 0.029
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Share Free-State peers (sd) of appointed state is the proportion of free companions calculated based on the place of appointment of graduates. Columns (1)-(3) focus on cadets from Slave States and Columns (4)-(6) on those from Free States. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
B.10 Correlation between Academic Rank and Military Rank in 1865
Table S13: Academic rank to military rank in 1865 (The Confederacy vs. The Union)
Dependent var Rank in 1865: war participants
The Confederacy The Union
(1) (2) (3) (4) (5) (6)
Class rank $0$0.001 -0.001 0.006∗∗∗ 0.006∗∗∗ 0.006∗∗
(0.003) (0.003) (0.003) (0.002) (0.002) (0.003)
Age in 1860 -0.004 -0.017 -0.034 -0.039
(0.063) (0.068) (0.043) (0.045)
Slave pop. share (sd) 0.148 0.029
(0.089) (0.063)
Cohort -0.049 -0.065 -0.087∗∗ -0.092∗∗
(0.061) (0.066) (0.042) (0.044)
State FEs N N Y N N Y
Dependent var.mean 7.516 7.516 7.516 5.848 5.848 5.848
Observations 281 281 281 637 637 637
R-squared 0.000 0.085 0.156 0.010 0.108 0.127
Note. The sample for this analysis consists of individuals who served in the Confederate Army (Columns (1)-(3)) and those who served in the Union Army (Columns (4)-(6)). This table shows that the association between academic rank and military rank is clear in the Union but not in the Confederacy. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.